OpenStudy (anonymous):

Does the table represent an exponential function? Explain. x:1:2:3:4: y:1:5:25:125 Ok, I would like to know what an exponential function is and how I can tell if one is or is not.

OpenStudy (anonymous):

yes, \[y=5^{x-1}\] when x=1, \[y=5^{1-1}=5^{0}=1\] when x=2 \[y=5^{2-1}=5\] actually it is a G.P. common ratio for y =5/1=5 first term a=1 \[tn=ar ^{n-1}\] Here tn=y,a=1,n=x

OpenStudy (anonymous):

r=5/1=5

OpenStudy (anonymous):

But I thought it had to be in the form?\[y=a(b)^3\]

OpenStudy (anonymous):

how you will get 1,5,25 ? it is best illustrated by \[y=5^{x-1},x=1,2,3,..\]